2f(x)+f(-x)=1/xsin(x-1/x).  find ∫f(x)dx in integral limits[e,1/e].

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sai

2f(x)+f(-x)=1/xsin(x-1/x) .............1

replace x by -x

2f(-x)+f(x)=1/xsin(x-1/x).............2

2*(1) -(2)

 f(x) = (1/3x) sin(x-1/x)

s0   I =limit [e,1/e]  ∫f(x)dx

         =limit [e,1/e]   ∫(1/3x) sin(x-1/x)  dx ................3

nor let x=1/t

  dx=  -1/t2  dt

 I=limit [1/e,e]   ∫(1/3t) sin(t-1/t)  dx

or I=limit [e,1/e]   -∫(1/3x) sin(x-1/x)  dx ..............4

add 3 and 4

2I =0

I=0

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the

answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin